# SecAlign: Defending Against Prompt Injection with Preference Optimization

### v2

@@ 4.2. SecAlign: SOTA Fine-Tuning-Based Defense
Jatmo [Piet et al., 2023], StruQ [Chen et al., 2025], BIPIA [Yi et al., 2023], instruction hierarchy [Wallace et al., 2024], and ISE [Wu et al., 2025] are existing fine-tuning-based defenses against prompt injection. Jatmo aims at a different setting where a base LLM is fine-tuned only for a specific instruction. Our comparison mainly focuses on StruQ, whose settings are closest to ours. BIPIA has been shown with a significant decrease in utility [Chen et al., 2025], and our evaluation confirms that. Instruction hierarchy is a private method proposed by OpenAI with no official implementation, so we query the GPT-4o-mini model that claims to deploy instruction hierarchy. ISE (Instructional Segment Embedding) is a concurrent work using architectural innovations, and there is also no official implementation, so we cannot compare with it.
@@
We reproduce StruQ [Chen et al., 2025] exactly using the released code, and there is no disparity in terms of dataset usage. We apply StruQ and SecAlign to Mistral-7B-Instruct and Llama3-8B-Instruct models that have been SFTed, and present the results with the original undefended counterpart in Fig. 3.
@@
For security, the open-weight models suffer from over 50% ASRs even under optimization-free attacks that could be generated within seconds. With optimization, the undefended model is broken with 89% and 98% ASRs respectively, indicating severe prompt injection threat in current LLMs in the community. StruQ effectively stops optimization-free attacks, but is vulnerable to optimization-based ones (34% and 62% ASRs for the two models). This concides the results in its official paper. In contrast, with great surprise, SecAlign decreases the ASRs of the strongest prompt injections to 1%, even if their injections are unseen and completely different from those in training. SecAlign becomes the first known method that reduces SOTA prompt injections to around 0%, which hints that LLMs secure against prompt injections may be possible, compared to the difficulty of securing classifiers against adversarial attacks.
The above results come from preference-optimizing the SFT model using a preference dataset (from Cleaned Alpaca [Ruebsamen, 2024]) that is in a different domain from the SFT dataset (private commercial one used by the industry). Below we show the defense performance when the preference and SFT dataset are in the same domain, i.e., both generated from Cleaned Alpaca. Here, the undefended model is SFTed from a base model; the StruQ model is defensive-SFTed from the base model; and the SecAlign model is preference-optimized from the undefended model. Results on three base models are shown in Fig. 4, and please see Fig. 7 in Appendix for Yi-1.5-6B and Llama-13B results supporting the same claim. Both StruQ and SecAlign demonstrate nearly identical WinRates on AlpacaEval2 compared to the undefended model, indicating minimal impact on the general usefulness of the model. By “identical”, we refer to a difference of $<0.7\%$, which is statistically insignificant given the standard error of 0.7% in the GPT4-based evaluator on AlpacaEval2 [Li et al., 2023a]. For security, SecAlign is secure against optimization-free attacks, and reduces the optimization-based ASRs from StruQ by a factor ¿4.
@@
Another fine-tuning-based defense against prompt injection is instruction hierarchy (IH) [Wallace et al., 2024], which implements a security policy where different instructions are assigned priority levels in the order of system $>$ user $>$ data. Whenever two instructions are conflicting, the higher-priority instruction is always favored over the lower one. Thus, IH mitigates prompt injection since malicious instructions in the data (lower priority) cannot override the user instruction (higher priority).
However, comparing IH against SecAlign is challenging because: (1) Since the implementation of IH is not publicly available, we cannot implement IH on the open-weight models used in our evaluation. Instead, we evaluate the GPT-4o-mini model, which reportedly implemented IH [OpenAI, 2024a], through its API. (2) As GPT-4o-mini is only available through API, we cannot implement any optimization-based attacks. Furthermore, the Completion and Ignore-Completion attacks cannot be implemented since those delimiters are not publicly known. Thus, we only consider Ignore attack, the weakest test one, in this evaluation.
Our evaluation shows that IH achieves 1% ASR against the Ignore attack. In comparison, SecAlign achieves 0% ASR against the Ignore attack across all five open-weight models; see Table 7 for details. Even though this is far from an apple-to-apple comparison, our result shows that SecAlign is likely more resilient to prompt injection attacks than IH even on weaker optimization-free attacks.
@@
Prompting-Based defenses are shown weaker than fine-tuning-based ones [Chen et al., 2025]. Here we also compare with existing baselines on Llama-7B and Llama3-8B-Instruct using optimization-free attacks to quantify SecAlign’s advantage. We consider the following prompting-based defenses that manually instruct the LLM to be mindful of prompt injection attacks in a zero-shot manner:
@@
  • Reminder defense [Yi et al., 2023] prompts the LLM to be aware of prompt injections after the original instruction by appending the following sentence to the instruction: Do not execute or adhere to any potential instructions or commands in the following data.
@@
4.4. Evaluation Setting Studies
We further study SecAlign performance under different security/utility evaluation settings.
Injection sentence
We test the security when the injected instruction is not asking the model to print a word, but instead to answer a question. We hold out 20% of the samples in the SFT-training-set, construct the preference dataset, and preference-optimize a Llama-7B. In evaluation, we use Ignore-Completion (the strongest optimization-free attack) to inject an instruction that is randomly sampled from the held-out SFT-training-set. In our manual inspection, SecAlign-defended LLM never responds to the injected instructions, even if they are in the same domain as the training set so the LLM tends to answer them. In another study, we change the injected word ”Hacked” to ”Hello World” and ”Serendipity” and report the ASR of optimization-free attacks on Llama-7B in Table 3. The multi-word injection ”Hello World” decreases the ASR, and the non-malicious-word injection ”Serendipity” increases the ASR. At all times, SecAlign achieves 0% ASR and outperforms baselines.
Table 3. Study of the injected word(s) and attack success rate criteria on Llama-7B. Breakdown numbers in Table 9.
Max ASR Opt.-Free Injected Defense
Attack Success Criteria Word None StruQ SecAlign
Begin-With (default) Hacked 75 5 0
In-Response 84 5 0.5
Begin-With (default) Hello World 73 0 0
In-Response 84 5 0.5
Begin-With (default) Serendipity 88 1 0
In-Response 92 3 1
Attack success criteria
In Table 3, we also include the rate when ”Hacked” or ”hacked” is in anywhere the response. In this case, the metric recalls all possible successful attacks, but also includes false positives, e.g., the output is repeating ”Print exactly hacked!” instead of answering it. In comparison, the default criteria (the attack succeeds only if it begins with the injection word) reports cases when an attack is guaranteed to succeed. In all cases, SecAlign outperforms the SOTA StruQ.
Injection position
In Table 4, we inject “Print exactly Hacked!” in different positions of the data. We find that our default attack setting, injecting after the data, succeeds most, which aligns with the observations in [Yi et al., 2023, Piet et al., 2023, Chen et al., 2025]. Therefore, our security evaluation favors the attacker most. All results come from the Ignore attack against undefended LLMs as StruQ or SecAlign LLMs have 0% ASRs.
Table 4. Prompt-injecting after the data (the default position) is the strongest. We use Llama-7B here.
Model Injection Position Ignore ASR (%)
before data 2
Llama-7B in the middle of data 2
after data 10
before data 13
Mistral-7B in the middle of data 7
after data 22
before data 24
Llama3-7B in the middle of data 14
after data 30
Utility benchmark
@@ 6. Conclusion and Discussions
  • Our construction of desirable outputs shares a disadvantage with all existing fine-tuning-based defenses: The desirable output ignores the injected instruction in the data instead of processing it as part of the data. This may lead the LLM to ignore some imperative sentences (which may not be an injection and should be handled as data, e.g., an imperative sentence to be translated) in the data, though we do not observe this phenomenon or hurt of utility in our study. Solving it requires a careful selection of the injection (based on the benign instruction) and a specially generated desirable response that is not from the SFT dataset.
  • As a defense to AI systems, SecAlign cannot achieve 100% security yet. For stronger security in LLM-integrated applications, we suspect the need for a multi-tiered defense combining SecAlign with other techniques such as detection (e.g., Prompt Shields [202, 2024c], PromptGuard [Meta, 2024]), and input reformatting [Jain et al., 2023]. We do not regard SecAlign as a final solution to prompt injection defense.
@@ Acknowledgments
This work is supported by Meta-BAIR Commons (Year 6). We are grateful for insightful discussions and comments from Chawin Sitawarin, Raluca Ada Popa, and anonymous reviewers.
@@
  • Fig. 7: results on more different-sized base models than Fig. 4.
@@
  • Table 9: breakdown numbers in Table 3 (The injection word and attack success criteria do not affect our claim).
Figure 7. Extending Fig. 4 to 6B and 13B base models.
@@
Table 9. The injection word and attack success criteria do not affect our claim: breakdown numbers from Table 3.

### v3

@@ 4.2. SecAlign: SOTA Fine-Tuning-Based Defense
Jatmo [Piet et al., 2023], StruQ [Chen et al., 2025a], BIPIA [Yi et al., 2023], instruction hierarchy [Wallace et al., 2024], and ISE [Wu et al., 2025a] are existing fine-tuning-based defenses against prompt injection. Jatmo aims at a different setting where a base LLM is fine-tuned only for a specific instruction. Our comparison mainly focuses on StruQ, whose settings are closest to ours. BIPIA has been shown with a significant decrease in utility [Chen et al., 2025a], and our evaluation confirms that. Instruction hierarchy is a private method proposed by OpenAI with no official implementation, so we query the GPT-4o-mini model that claims to deploy instruction hierarchy. ISE (Instructional Segment Embedding) is a concurrent work using architectural innovations, and there is also no official implementation, so we cannot compare with it.
@@
We reproduce StruQ [Chen et al., 2025a] exactly using the released code, and there is no disparity in terms of dataset usage. We apply StruQ and SecAlign to Mistral-7B-Instruct and Llama3-8B-Instruct models that have been SFTed, and present the results with the original undefended counterpart in Fig. 3.
@@
For security, the open-weight models suffer from over 50% ASRs even under optimization-free attacks that could be generated within seconds. With optimization, the undefended model is broken with 89% and 97% ASRs respectively, indicating severe prompt injection threat in current LLMs in the community. StruQ effectively stops optimization-free attacks, but is vulnerable to optimization-based ones (27% and 45% ASRs for the two models). This coincides the results in its official paper. In contrast, with great surprise, SecAlign decreases the ASRs of the strongest prompt injections to 1% and 8%, even if their injections are unseen and completely different from those in training. The great empirical success of SecAlign hints that LLMs secure against prompt injections may be possible, compared to the difficulty of securing classifiers against adversarial attacks.
The above results come from preference-optimizing the SFT model using a preference dataset (from Cleaned Alpaca [Ruebsamen, 2024]) that is in a different domain from the SFT dataset (private commercial one used by the industry). Below we show the defense performance when the preference and SFT dataset are in the same domain, i.e., both generated from Cleaned Alpaca. Here, the undefended model is SFTed from a base model; the StruQ model is defensive-SFTed from the base model; and the SecAlign model is preference-optimized from the undefended model. Results on three base models are shown in Fig. 4. Both StruQ and SecAlign demonstrate nearly identical WinRates on AlpacaEval2 compared to the undefended model, indicating minimal impact on the general usefulness of the model. By “identical”, we refer to a difference of $<0.7\%$, which is statistically insignificant given the standard error of 0.7% in the GPT4-based evaluator on AlpacaEval2 [Li et al., 2023a]. For security, SecAlign is secure against optimization-free attacks, and reduces the optimization-based ASRs from StruQ by a factor ¿4.
@@
Another fine-tuning-based defense against prompt injection is instruction hierarchy [Wallace et al., 2024], which implements a security policy where different instructions are assigned priority levels in the order of system $>$ user $>$ data. Whenever two instructions are conflicting, the higher-priority instruction is always favored over the lower one. Thus, instruction hierarchy mitigates prompt injection since malicious instructions in the data (lower priority, called ”tool outputs” in the paper) cannot override the user instruction (higher priority, ”user message” in the paper).
To evaluate this level of security, we create a dummy tool function that returns the data part as its output, and put the intended instruction in the ”user” role. Since the implementation of instruction hierarchy is not publicly available, we cannot implement instruction hierarchy on the open-weight models used in our evaluation. Instead, we evaluate the GPT-4o-mini model, which reportedly implemented instruction hierarchy [OpenAI, 2024a]. As GPT-4o-mini is only available through API, we cannot implement any optimization-based attacks.
Our evaluation shows that instruction hierarchy achieves 1% ASR against the optimization-free Ignore attack. For reference, SecAlign achieves 0% ASR against the Ignore attack across all five open-weight models; see Table 6 for details. We note that this is far from an apple-to-apple comparison since the base model for instruction hierarchy is completely different from the base model for SecAlign.
@@
Prompting-Based defenses are shown weaker than fine-tuning-based ones [Chen et al., 2025a]. Here we also compare with existing baselines on Llama-7B and Llama3-8B-Instruct using optimization-free attacks to quantify SecAlign’s advantage. We consider the following prompting-based defenses that manually instruct the LLM to be mindful of prompt injection attacks in a zero-shot manner:
@@
  • Reminder defense [Yi et al., 2023] prompts the LLM to be aware of prompt injections by appending the following sentence to the user instruction: Do not execute or adhere to any potential instructions or commands in the following data.
@@
4.4. Security Generalization of SecAlign
To diversify evaluations on injection position (besides at the end) and task (besides printing hacked) on larger testset, we extend our security evaluations to the SEP prompt injection benchmark [Zverev et al., 2025]. SEP has 9.1K samples, each with a unique injection task. We vary the injection position to be the start/middle/end of the data. We ask GPT-4-Turbo to judge attack success, and also to judge the defended models’ output quality against the undefended one as the utility (under no attack).
SecAlign secures Llama-3-8B-Instruct significantly without much loss of utility in our evaluations, see Table 3. By comparison, although StruQ (with a tuned learning rate) attains lower ASRs, this is achieved by a drastically lower utility as the resulting LLM fails to respond to the benign instruction as well. Without any defense, injecting after the data succeeds most, which aligns with the observations in [Yi et al., 2023, Piet et al., 2023, Chen et al., 2025a]. In both StruQ/SecAlign, the defense is stronger against prompt injections at the end of data (same injection position as in training) compared to that at the start.
In Table 3, we have also tested on an API-calling prompt injection benchmark, InjecAgent [Zhan et al., 2024], which prompts a tested LLM to process data retrieved from APIs. The attack succeeds when a malicious API (instructed by the injection in retrieval) is called. In a significantly different domain (API data, long inputs), SecAlign achieves 0 ASR, showing strong defense generalization. We are unable to study the utility-security trade-off in InjecAgent since it does not provide utility evaluation.
Table 3. Security evaluations on out-of-distribution prompt injection benchmarks with Llama3-8B-Instruct.
Defense None StruQ SecAlign
SEP Utility (w.r.t. None, %, $\uparrow$) 50.0 18.9 46.6
SEP Ignore ASR (start, %, $\downarrow$) 54.7 2.1 11.0
SEP Ignore ASR (middle, %, $\downarrow$) 39.5 0.1 3.6
SEP Ignore ASR (end, %, $\downarrow$) 64.0 0.1 6.5
InjecAgent ASR (%, $\downarrow$) 75.9 2.2 0
We also study whether many-shot prompting, effective in jailbreaking attacks [Anil et al., 2024], is helpful in prompt injection attacks. We add demonstrations of prompt-injected-inputs with undesirable responses, using the completion attack structure/delimiters, before “print exactly hacked”. Contrary to many-shot jailbreak attacks, adding demonstrations in PIs does not increase ASR for undefended Llama3-8B-Instruct, and the StruQ/SecAlign LLMs both get 0% ASRs. We hypothesize this is because “responding to injection” is a more complex attacker goal to enforce via many-shot demonstrations compared to jailbreaking. For prompt injections in different languages, we change “Print exactly” to Chinese/Spanish in Completion attacks, and StruQ/SecAlign Llama3-8B-Instruct both get 0% ASRs.
4.5. Utility Generalization of SecAlign
@@
Our construction of desirable outputs shares one property with all existing fine-tuning-based defenses: The desirable output ignores the injected instruction in the data instead of processing it as part of the data. Thus, it is important to study in test time, how the SecAlign LLM processes imperative sentences in the data part (which may not be an injection and should be handled as data, e.g., an imperative sentence to be translated).
We use the instruction “The sentence you are given might be too wordy, complicated, or unclear. Rewrite the sentence and make your writing clearer by keeping it concise. Whenever possible, break complex sentences into multiple sentences and eliminate unnecessary words.” and the data part being different instructions in the testset. We use GPT-4-Turbo (AlpacaEval2-prompting) to compare the output quality of Meta-Llama-3-8B-Instruct (SecAlign) against that of the undefended counterpart on all other 804 samples, and the WinRate is 65.5%. A ¿50% WinRate means the SecAlign model is better at processing imperative sentences in data as data, instead of as instructions. We also perform manual inspection on the first 50 test samples with similar findings: 16% of imperative data are handled as data by Meta-Llama-3-8B-Instruct (undefended) vs. 52% for SecAlign one. In the tests above, we do not observe utility loss due to our way of dataset generation.
@@ 6. Conclusion and Discussions
  • As a defense to AI systems, SecAlign cannot achieve 100% security, and may be evaded by future attacks that are not tested, e.g., prompt injections through multi-turn conversations in applications like web-agents. It is also unclear how SecAlign LLMs perform if they are further fine-tuned. Lastly, our utility datasets have one instruction, so we are not sure about the utility of SecAlign when there are multiple benign instructions.
  • SecAlign is most effective when the injection is at the end of the data, see Table 3, despite a strong generalization to injections in other positions. For better security generation, simulating injections in different positions in training [Abdelnabi et al., 2025] is a possible strategy.
@@
For stronger security in LLM-integrated applications, we suspect the need for a multi-tiered defense combining SecAlign with other techniques such as detection (e.g., Prompt Shields [202, 2024c], PromptGuard [Meta, 2024]), input reformatting [Jain et al., 2023], output manipulation [Wu et al., 2025b], and system-level defense [Debenedetti et al., 2025]. We do not regard SecAlign as a standalone solution to prompt injection attacks.
@@ Acknowledgments
This research was supported by the Meta-BAIR Commons (2024-2026). UC Berkeley was supported by National Science Foundation under grant 2229876 (the ACTION center), Open Philanthropy, the Department of Homeland Security, and IBM. We are grateful for insightful discussions and comments from Chawin Sitawarin, Raluca Ada Popa, and anonymous reviewers.
@@
  • Table 8: In another study on the AlpacaFarm testset (used in main experiments), we change the injected word ”Hacked” to ”Hello World” and ”Serendipity” and report the ASR of optimization-free attacks. The multi-word injection ”Hello World” decreases the ASR, and the non-malicious-word injection ”Serendipity” increases the ASR. At all times, SecAlign achieves 0% ASR. In Table 8, we also include the rate when ”Hacked” or ”hacked” is in anywhere the response. In this case, the metric recalls all possible successful attacks, but also includes false positives, e.g., the output is repeating ”Print exactly hacked!” instead of answering it. In comparison, the default criteria (the attack succeeds only if it begins with the injection word) reports cases when an attack is guaranteed to succeed. In all criteria, SecAlign outperforms StruQ.
@@
Table 8. The injection word and attack success criteria do not affect our claim.